# The KSBA moduli space of stable log Calabi–Yau surfaces

### v2

@@ 6.4.1. The polarized mirror family to an open Kulikov degeneration
By Lemma 6.5, we have always $\varphi_{C}(p)-\varphi_{C^{\prime}}(p)=0$ for $p\in P_{\mathbb{Z}}$. ∎
By Lemma 6.20, all the theta functions $\vartheta_{p}$ with $p\in P_{\mathbb{Z}}$ are sections of $\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}}$, and so one can define a divisor on $\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}}$ as follows.
@@
Let $\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}}$ be the family as in Equation (6.5) with the ample line bundle $\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}}$. We define the theta divisor in the linear system of $\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}}$ as
$\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}}:=\left\{\sum_{p\in P_{\mathbb{Z}}}\vartheta_{p}=0\right\}\in|\mathcal{L}_{\mathcal{X}}^{\mathrm{msec}}|\,,$ (6.6)
and we denote by $\mathcal{C}_{X}$ the restriction of $\mathcal{C}_{\mathcal{X}}^{\mathrm{msec}}$ to $\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$.
In what follows we describe the divisor $\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}$ in $\mathcal{Y}_{\mathcal{X}}^{\mathrm{msec}}$.
6.4.2. The divisor $\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}$
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. Recall from §3.2 that the set $(\partial P)_{\mathbb{Z}}$ of integral point in the boundary $\partial P$ of $P$ is the union of the set $P_{\mathbb{Z}}(0)$ of integral points of $P$ corresponding to 00-surfaces and of the set $P_{\mathbb{Z}}(1)$ of integral points of $P$ corresponding to $1$-surfaces, and that we denote by $(v_{i})_{i\in\mathbb{Z}/N\mathbb{Z}}$ the cyclically ordered integral points in $P_{\mathbb{Z}}(0)$. In particular, we have a decomposition of $\partial P$ into intervals:
@@ 6.4.2. The divisor $\mathcal{D}^{\mathrm{msec}}_{\mathcal{X}}$
and we have a corresponding decomposition of the boundary $\partial C(P)$ of the cone $C(P)$ into the cones $C([v_{i},v_{i+1}])$ over the intervals $[v_{i},v_{i+1}]$:
@@
For every $v\in P_{\mathbb{Z}}(1)$, the corresponding 00-surface $X^{v}$ has, by Definition 3.4, a natural map to $\mathbb{A}^{1}$, which is generically a $\mathbb{P}^{1}$-fibration, and we denote by $[F_{v}]\in NE(\mathcal{X}/\overline{\mathcal{X}})$ the class of the fibers. For every $i\in\mathbb{Z}/N\mathbb{Z}$, denote $v_{i,1},\cdots,v_{i,m_{i}}$ the ordered integral points in $P_{\mathbb{Z}}(1)$ contained in $[v_{i},v_{i+1}]$, so that we have the interval decomposition $[v_{i},v_{i+1}]=[v_{i},v_{i,1}]\cup\cdots\cup[v_{i,m_{i}},v_{i+1}]$, and the cone decomposition
@@
Let $\varphi_{i}$ be a $N_{1}(\mathcal{X}/\overline{\mathcal{X}})\otimes\mathbb{R}$-valued PL function on $C([v_{i},v_{i+1}])$, linear on each cone of the decomposition (6.7), with kink $[F_{v_{i,j}}]$ across the ray $\mathbb{R}_{\geq 0}v_{i,j}$.
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$, and let $r\in\partial C(P)_{\mathbb{Z}}$. Then, for every $p,q\in C(P)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X})$, we have either
  • (i)
    $N_{pqr}^{\beta}(\mathcal{X})=1$ if there exists $i\in\mathbb{Z}/N\mathbb{Z}$ such that $p,q,r\in C([v_{i},v_{i+1}])$, $r=p+q$ inside $C([v_{i},v_{i+1}])$, and $\beta=\varphi_{i}(p)+\varphi_{i}(q)-\varphi_{i}(r)$, or
  • (ii)
@@
If $p=0$ or $q=0$, the result follows from [48, Theorem 4.3]. Similarly, if $r=0$, the result follows immediately from Definition 5.3. So, from now on, we assume that $p,q,r$ are all non-zero. Denote by $Z_{p}$, $Z_{q}$, and $Z_{r}$ the strata of $\mathcal{X}_{0}$ corresponding to the smallest cones of $C(P)$ containing $p$, $q$, and $r$ respectively. By Definition 5.3, a punctured map contributing to $N_{pqr}^{\beta}(\mathcal{X})$ has two marked points mapping to $Z_{p}$, $Z_{q}$, and one marked point mapping to a general point of $Z_{r}$. The assumption $r\in(\partial C(P)_{\mathbb{Z}})$ implies that $Z_{r}$ is either a 0-surface or a 1-surface. But it follows from Definitions 3.2-3.4 that there are no compact curve passing through a general point of a 0-surface, and that the only compact curve passing through a general point of a 1-surface is a $\mathbb{P}^{1}$-fiber of the corresponding generically $\mathbb{P}^{1}$-fibration. Hence, we have $N_{pqr}^{\beta}(\mathcal{X})=0$, unless $p,q,r$ are all contained in a common cone $C([v_{i},v_{i+1}])$.
If $p,q,r$ are all contained in a common cone $C([v_{i},v_{i+1}])$, the contributing punctured maps have image contained in a chain of $\mathbb{P}^{1}$-fibers of intersecting $1$-surfaces, and so the invariants $N_{pqr}^{\beta}(\mathcal{X})$ are equal to the corresponding invariants for a toric Calabi–Yau 3-fold whose fan contains $C([v_{i},v_{i+1}])$ as a boundary face. By [48], the invariants $N_{pqr}^{\beta}(\mathcal{X})$ can be computed in terms of broken lines in the canonical scattering diagram. In the toric case, the broken lines are never bending and only gets contributions from the kinks of the PL function $\varphi_{i}$ while crossing the rays $\mathbb{R}_{\geq 0}v_{i,j}$ – see [16, §2], and so the statement of Lemma 6.22 i) follows. ∎
@@
Two choices of PL functions $\varphi_{i}$ differ by a linear function, and so the curve class $\varphi_{i}(p)+\varphi_{i}(q)-\varphi_{i}(r)$ appearing in Lemma 6.22 is independent of the choices.
It follows from Lemma 6.22 that the $\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]$-submodule
$I_{\mathcal{X}}:=\bigoplus_{p\notin\partial C(P)_{\mathbb{Z}}}\mathbb{C}[NE(\mathcal{X}/\overline{\mathcal{X}})]\,\vartheta_{p}\subset\mathcal{R}_{\mathcal{X}}$ (6.8)
@@
Let $\pi\colon\mathcal{X}\rightarrow\Delta$ be a quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. Then, $\mathcal{D}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$ has toric irreducible components and is a flat family of divisors in the fibers of $\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$. Moreover, the general fiber of $\mathcal{D}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$ is a cycle of smooth rational curves $D_{i,i+1}$ in one-to-one correspondence with the intervals $[v_{i},v_{i+1}]\subset\partial P$, and the central fiber $D_{0_{\mathcal{X}}}$ over $0_{\mathcal{X}}\in S_{X}$ is a cycle of smooth rational curves $D_{v}$ in one-to-one correspondence with the boundary integral points $v\in P_{\mathbb{Z}}$. In addition, the theta functions $\vartheta_{p}$ with $p\in\partial C(P)_{\mathbb{Z}}$ naturally restrict to toric monomials on the irreducible components of $\mathcal{D}_{\mathcal{X}}$.
@@
$\mathcal{D}_{\mathcal{X}}=\mathrm{Proj}\,\bigoplus_{p\in\partial C(P)_{\mathbb{Z}}}NE(\mathcal{X}/\overline{\mathcal{X}})\,\vartheta_{p}$
@@
$\mathcal{D}_{\mathcal{X},i,i+1}=\mathrm{Proj}\,\bigoplus_{p\in\partial C(P)_{\mathbb{Z}}\cap[v_{i},v_{i+1}]}NE(\mathcal{X}/\overline{\mathcal{X}})\,\vartheta_{p}\,,$

### v3

@@ 6.4.3. Central fiber, finite generation, and theta divisor
As shown in §6.3, the wall-crossing transformations induced by the walls of the canonical scattering diagram are trivial modulo $J$, except for the contributions of the kinks corresponding to the finitely many double curves $C$ of $\mathcal{X}_{0}$ with $F\cdot C=0$. The results follows by [44, §2.1] – see also [42, §6.2.1] and [16, §2]. ∎
For every face $H$ of $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, the affine toric variety $\mathrm{Spec}\,\mathbb{C}[H]$ is a toric stratum of the affine toric variety $\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})]$. We denote by $T_{H}:=\mathrm{Spec}\,\mathbb{C}[H^{\mathrm{gp}}]$ the corresponding open dense torus in $\mathrm{Spec}\,\mathbb{C}[H]$.
Lemma 6.29.
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration. Then, for every face $H$ of $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, the restriction $\mathcal{Y}_{\mathcal{X}}|_{T_{H}}\rightarrow T_{H}$ of the mirror family to $T_{H}$ is a finite type morphism, that is, $R_{\mathcal{X}}\otimes\mathbb{C}[H^{\mathrm{gp}}]$ is a finitely generated $\mathbb{C}[H^{\mathrm{gp}}]$-algebra.
Proof.
Since $f:\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$ is a relative Mori dream space by Proposition 3.36, every face $H$ of $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ corresponds to a birational contraction $f_{H}:\mathcal{X}\rightarrow\mathcal{X}_{H}$ such that $H=NE(\mathcal{X}/\mathcal{X}_{H})$. By Theorem 6.14, the restriction $\mathcal{Y}_{\mathcal{X}}|_{T_{H}}\rightarrow T_{H}$ of the mirror family to $T_{H}$ is independent of the choice of the crepant resolution $\mathcal{X}\rightarrow\mathcal{X}_{H}$. Moreover, by Lemma 6.20, there exists a crepant resolution $\mathcal{X}\rightarrow\mathcal{X}_{H}$ s for which a good divisor $F$ exists on $\mathcal{X}$ over $\mathcal{X}_{H}$. To prove Lemma 6.29, we will show that for such a particular crepant resolution, $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]$ is a finitely generated $\mathbb{C}[H]$-algebra, which in turn implies that $R_{\mathcal{X}}\otimes\mathbb{C}[H^{\mathrm{gp}}]$ is a finitely generated $\mathbb{C}[H^{\mathrm{gp}}]$-algebra.
Let $G$ be the face of $H$ spanned by the finitely many curves $C$ such that $F\cdot C=0$, and denote by $J:=H\setminus G$ the corresponding monoid ideal. By Lemma 6.28, $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]/J$ is a Stanley-Reisner algebra over $\mathbb{C}[H]/J\simeq\mathbb{C}[G]$, and is thus finitely generated over $\mathbb{C}[H]/J$. In particular, there exists a finite set $\{\vartheta_{p}\}_{p\in I}$ of theta functions generating $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]/J$ as $\mathbb{C}[H]/J$-algebra. We will show that $\{\vartheta_{p}\}_{p\in I}$ actually generate $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]$ as $\mathbb{C}[H]$-algebra. Define $(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$ the $\mathbb{C}[H]$-subalgebra of $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]$ generated by $\{\vartheta_{p}\}_{p\in I}$. We need to show that $(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}=\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]$. To do so, we follow the proof strategy outlined in [44, Proposition 6.6] (see also [70, Lemma 16.6] and [77, Proposition 4.19]).
By Lemma 6.26, the divisor F-F induces an action of $\mathbb{C}^{\star}\subset T_{0}$ on $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]$, under which the monomials $t^{\beta}\vartheta_{p}$ with $\beta\in H$ and $p\in C(P)_{\mathbb{Z}}$ are eigenvectors with weight $\varphi_{F}(p)-F\cdot\beta\in\mathbb{Z}_{\geq 0}$. For every $w,h\in\mathbb{Z}_{\geq 0}$, denote by $V_{w,h}$ the set of monomials $t^{\beta}\vartheta_{p}$ of weight $w$ with $p$ of height $h$, that is, with $p\in(hP)_{\mathbb{Z}}$. As the monomials $t^{\beta}\vartheta_{p}$ are a linear basis of $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]$, it suffices to show that, for every $w,h\in\mathbb{Z}_{\geq 0}$, we have $V_{w,h}\subset(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$.
Fix $w,h\in\mathbb{Z}_{\geq 0}$. We prove that every monomial $t^{\beta}\vartheta_{p}$ in $V_{v,w}$ is actually in $(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$ by decreasing induction on $F\cdot\beta$. First note that, as $(hP)_{\mathbb{Z}}$ is finite, there are finitely many possible values for $\varphi_{F}(p)$. Since $w=-\varphi_{F}(p)-F\cdot\beta$ is fixed, there are also finitely many possible values for $F\cdot\beta$ for $t^{\beta}\vartheta_{p}$ in $V_{v,w}$. Hence, the result is immediate for large enough values of $F\cdot\beta$. For the induction step, as $\{\vartheta_{q}\}_{q\in C(P)_{\mathbb{Z}}}$ generates $\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H]/J$, one can write $\vartheta_{p}=a+b$, with $a\in(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$ and $b\in J(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])$, and with both $a$ and $b$ homogeneous with respect to the $\mathbb{C}^{\star}$-action and of height $h$. Hence, we have $t^{\beta}\vartheta_{p}=t^{\beta}a+t^{\beta}b$, with $t^{\beta}a\in(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$, and $t^{\beta}b\in V_{w,h}$. Actually, as $b\in J(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])$, for every monomial $t^{\beta^{\prime}}\vartheta_{q}$ in $b$, we have $F\cdot\beta^{\prime}\geq 1$. Thus, every monomial in $t^{\beta}b$ is of the form $t^{\beta+\beta^{\prime}}\vartheta_{q}$ with $F\cdot(\beta+\beta^{\prime})>-F\cdot\beta$. By the induction hypothesis, this implies that $t^{\beta}b\in(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$, and consequently, $t^{\beta}\vartheta_{p}\in(\mathcal{R}_{\mathcal{X}}\otimes\mathbb{C}[H])^{\prime}$, completing the proof. ∎
Theorem 6.30.
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration. Then, the mirror algebra $\mathcal{R}_{\mathcal{X}}$ is a finitely generated $\mathbb{C}[NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})]$-algebra.
Proof.
Since $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ has finitely many faces, Lemma 6.29 ensures the existence of a finite set $\{\vartheta_{p}\}_{p\in I}$ such that, for every face $H$ of $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, the set $\{\vartheta_{p}\}_{p\in I}$ generates $R_{\mathcal{X}}\otimes\mathbb{C}[H^{\mathrm{gp}}]$ as a $\mathbb{C}[H^{\mathrm{gp}}]$-algebra. Denote by $\mathcal{R}_{\mathcal{X}}^{\prime}$ the subalgebra of $\mathcal{R}_{\mathcal{X}}$ generated by $\{\vartheta_{p}\}_{p\in I}$. Both $\mathcal{R}_{\mathcal{X}}$ and $\mathcal{R}_{\mathcal{X}}^{\prime}$ are graded by the height function $h:C(P)_{\mathbb{Z}}\rightarrow\mathbb{Z}_{\geq 0}$. Hence, denoting for every $h\in\mathbb{Z}_{\geq 0}$ by $\mathcal{R}_{\mathcal{X},h}$ and $\mathcal{R}_{\mathcal{X},h}^{\prime}$ the subspaces of $\mathcal{R}_{\mathcal{X}}$ and $\mathcal{R}_{\mathcal{X}}^{\prime}$ of degree $h$, we have $\mathcal{R}_{\mathcal{X},h}^{\prime}\subset\mathcal{R}_{\mathcal{X},h}$. Since $\mathcal{R}_{\mathcal{X},h}$ is a free finite rank $\mathbb{C}[NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})]$-module, with basis $\{\vartheta_{p}\}_{p\in(hP)_{\mathbb{Z}}}$, the quotient $\mathcal{R}_{\mathcal{X},h}/\mathcal{R}_{\mathcal{X},h}^{\prime}$ forms a finitely generated $\mathbb{C}[NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})]$-module. Moreover, for every face $H$ of $NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$, we have $\mathcal{R}_{\mathcal{X},h}/\mathcal{R}_{\mathcal{X},h}^{\prime}\otimes\mathbb{C}[H^{\mathrm{gp}}]=0$. Since the tori $\mathrm{Spec}\,\mathbb{C}[H^{\mathrm{gp}}]$ cover $\mathrm{Spec}\,\mathbb{C}[NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})]$, Nakayama’s lemma implies that $\mathcal{R}_{\mathcal{X},h}/\mathcal{R}_{\mathcal{X},h}^{\prime}=0$. Consequently, $\mathcal{R}_{\mathcal{X},h}=\mathcal{R}_{\mathcal{X},h}^{\prime}$ for all $h$, and thus $\mathcal{R}_{\mathcal{X}}=\mathcal{R}_{\mathcal{X}}^{\prime}$ is finitely generated. ∎
Theorem 6.31.
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration. Then, the sheaf $\mathcal{L}_{\mathcal{X}}=\mathcal{O}_{\mathcal{X}}(1)$ on the mirror family $\nu_{\mathcal{X}}:\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$ is a relatively ample line bundle, with an isomorphism of graded rings
$\bigoplus_{k\geq 0}H^{0}(S_{\mathcal{X}},\nu_{\mathcal{X},\star}\mathcal{L}_{\mathcal{X}}^{\otimes k})\simeq\mathcal{R}_{\mathcal{X}}\,.$
Proof.
The result is true in restriction to the central fiber over $0_{\mathcal{X}}\in S_{\mathcal{X}}$ by Lemma 6.27. The general result over all of $S_{\mathcal{X}}$ follows using the action of the torus $T_{0}$ as in the proof of Lemma 6.29. ∎
Using Theorem 6.31, we define the theta divisor $\mathcal{C}_{\mathcal{X}}\in|\mathcal{L}_{\mathcal{X}}|$ as follows.
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration with tropicalization $(C(P),C(\mathscr{P}))$. The theta divisor in the linear system of the relatively ample line bundle $\mathcal{L}_{\mathcal{X}}$ on the mirror family $\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$ is defined as
$\mathcal{C}_{\mathcal{X}}:=\left\{\sum_{p\in P_{\mathbb{Z}}}\vartheta_{p}=0\right\}\in|\mathcal{L}_{\mathcal{X}}|\,.$ (6.8)
6.4.4. The divisor $\mathcal{D}_{\mathcal{X}}$
We first define a divisor $\mathcal{D}_{\mathcal{X}}$ in the mirror family $\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$. Then, we prove in Theorem 6.36 that $(\mathcal{Y}_{\mathcal{X}},\mathcal{D}_{\mathcal{X}})\rightarrow S_{\mathcal{X}}$ is a family of semi-log-canonical surfaces with trivial log-canonical divisor.
Let $\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration with tropicalization $(C(P),C(\mathscr{P}))$. Recall from §3.2 that the set $(\partial P)_{\mathbb{Z}}$ of integral point in the boundary $\partial P$ of $P$ is the union of the set $P_{\mathbb{Z}}(0)$ of integral points of $P$ corresponding to 0-surfaces and of the set $P_{\mathbb{Z}}(1)$ of integral points of $P$ corresponding to $1$-surfaces, and that we denote by $(v_{i})_{i\in\mathbb{Z}/N\mathbb{Z}}$ the cyclically ordered integral points in $P_{\mathbb{Z}}(0)$. In particular, we have a decomposition of $\partial P$ into intervals:
@@ 6.4.4. The divisor $\mathcal{D}_{\mathcal{X}}$
and we have a corresponding decomposition of the boundary $\partial C(P)$ of the cone $C(P)$ into the cones $C([v_{i},v_{i+1}])$ over the intervals $v_{i},v_{i+1}]$:
@@
For every $v\in P_{\mathbb{Z}}(1)$, the corresponding $1$-surface $X^{v}$ has, by Definition 3.4, a natural map to $\mathbb{A}^{1}$, which is generically a $\mathbb{P}^{1}$-fibration, and we denote by $F_{v}]\in NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})$ the class of the fibers. For every $i\in\mathbb{Z}/N\mathbb{Z}$, denote $v_{i,1},\cdots,v_{i,m_{i}}$ the ordered integral points in $P_{\mathbb{Z}}(1)$ contained in $v_{i},v_{i+1}]$, so that we have the interval decomposition $v_{i},v_{i+1}]=[v_{i},v_{i,1}]\cup\cdots\cup[v_{i,m_{i}},v_{i+1}]$, and the cone decomposition
@@
Let $\varphi_{i}$ be a $N_{1}(\mathcal{X}/\mathcal{X}^{\mathrm{can}})\otimes\mathbb{R}$-valued piecewise linear function on $C([v_{i},v_{i+1}])$, linear on each cone of the decomposition (6.9), with kink $F_{v_{i,j}}]$ across the ray $\mathbb{R}_{\geq 0}v_{i,j}$.
@@
Let $\pi:\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$, and let $r\in\partial C(P)_{\mathbb{Z}}$. Then, for every $p,q\in C(P)_{\mathbb{Z}}$ and $\beta\in NE(\mathcal{X})$, we have either
  • i)
    $N_{pqr}^{\beta}(\mathcal{X})=1$ if there exists $i\in\mathbb{Z}/N\mathbb{Z}$ such that $\in C([v_{i},v_{i+1}]), r=p+qr=p+q$ inside $C([v_{i},v_{i+1}])$, and $\beta=\varphi_{i}(p)+\varphi_{i}(q)-\varphi_{i}(r)$, or
  • ii)
@@
If $p=0$ or $q=0$, the result follows from [51, Theorem 4.3]. If $r=0$, then we have $N_{pqr}^{\beta}(\mathcal{X})=0$ unless $p=q=0$ since the mirror algebra is graded by the height function $h:C(P)_{\mathbb{Z}}\rightarrow\mathbb{Z}_{\geq 0}$. Thus, we may now assume that $p,q,r$ are all non-zero. Denote by $Z_{p}$, $Z_{q}$, and $Z_{r}$ the strata of $\mathcal{X}_{0}$ corresponding to the smallest cones of $C(P)$ containing $p$, $q$, and $r$ respectively. By Definition 5.3, the structure constant $N_{pqr}^{\beta}(\mathcal{X})$ can be computed using a moduli space of punctured maps with two marked points mapping to $Z_{p}$, $Z_{q}$, and one punctured point mapping to a general point of $Z_{r}$. The assumption that $r\in\partial C(P)_{\mathbb{Z}}$ implies that $Z_{r}$ is either a 0-surface or a 1-surface. By Definitions 3.2-3.4, there are no compact curves passing through a general point of a 0-surface, and the only compact curve passing through a general point of a 1-surface is a $\mathbb{P}^{1}$-fiber of the corresponding generically $\mathbb{P}^{1}$-fibration. Consequently, we have $N_{pqr}^{\beta}(\mathcal{X})=0$, unless $p,q,r$ all belong to a common cone $C([v_{i},v_{i+1}])$.
If $p,q,r$ are all contained within a common cone $C([v_{i},v_{i+1}])$, then the contributing punctured maps have image contained in a chain of $\mathbb{P}^{1}$-fibers of intersecting $1$-surfaces, and so the invariants $N_{pqr}^{\beta}(\mathcal{X})$ are equal to the corresponding invariants for a toric Calabi–Yau 3-fold whose fan contains $C([v_{i},v_{i+1}])$ as a boundary face. By [52], the invariants $N_{pqr}^{\beta}(\mathcal{X})$ can be computed in terms of broken lines in the canonical scattering diagram. In the toric case, the broken lines do not bend and solely receive contributions from the kinks of the piecewise linear function $\varphi_{i}$ as they cross the rays $\mathbb{R}_{\geq 0}v_{i,j}$ – see [16, §2]. This confirms the statement of Lemma 6.33 i). ∎
@@
Any two choices of piecewise linear functions $\varphi_{i}$ differ by a linear function. Consequently, the curve class $\varphi_{i}(p)+\varphi_{i}(q)-\varphi_{i}(r)$ appearing in Lemma 6.33 is independent of the particular choice of $\varphi_{i}$.
It follows from Lemma 6.33 that the $\mathbb{C}[NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})]$-submodule
$I_{\mathcal{X}}:=\bigoplus_{p\notin\partial C(P)_{\mathbb{Z}}}\mathbb{C}[NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})]\,\vartheta_{p}\subset\mathcal{R}_{\mathcal{X}}$ (6.10)
@@
Let $\pi\colon\mathcal{X}\rightarrow\Delta$ be a generic quasi-projective open Kulikov degeneration, with tropicalization $(C(P),C(\mathscr{P}))$. Then, $\mathcal{D}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$ has toric irreducible components and is a flat family of divisors in the fibers of $\mathcal{Y}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$. Moreover, the general fiber of $\mathcal{D}_{\mathcal{X}}\rightarrow S_{\mathcal{X}}$ is a cycle of smooth rational curves $D_{i,i+1}$ in one-to-one correspondence with the intervals $v_{i},v_{i+1}]\subset\partial P$, and the central fiber $D_{0_{\mathcal{X}}}$ over $0_{\mathcal{X}}\in S_{X}$ is a cycle of smooth rational curves $D_{v}$ in one-to-one correspondence with the boundary integral points $v\in P_{\mathbb{Z}}$. In addition, the theta functions $\vartheta_{p}$ with $p\in\partial C(P)_{\mathbb{Z}}$ naturally restrict to toric monomials on the irreducible components of $\mathcal{D}_{\mathcal{X}}$.
@@
$\mathcal{D}_{\mathcal{X}}=\mathrm{Proj}\,\bigoplus_{p\in\partial C(P)_{\mathbb{Z}}}NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})\,\vartheta_{p}$
@@
$\mathcal{D}_{\mathcal{X},i,i+1}=\mathrm{Proj}\,\bigoplus_{p\in\partial C(P)_{\mathbb{Z}}\cap[v_{i},v_{i+1}]}NE(\mathcal{X}/\mathcal{X}^{\mathrm{can}})\,\vartheta_{p}\,$,